OpenStudy (anonymous):

Solve. show the steps to solve equation 2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

n = 2

OpenStudy (anonymous):

show how you got it please

OpenStudy (hoblos):

2(n-3) +5 = 3(n-1) => 2n-6+5=3n-3 => -n = -2 => n=2